OpenStudy (anonymous):

initial value problem -16cos(4t+pi), s'(0)=200, s(0)=0

OpenStudy (dumbcow):

shouldn't there be a s'(t) somewhere

OpenStudy (anonymous):

its actually d^2s/dt^2=-16cos(4t+pi), s'(0)=200, s(0)=0

OpenStudy (anonymous):

i am having trouble in find the initial value

OpenStudy (dumbcow):

ok so you need to integrate function to find ds/dt, then integrate again to find s(t)

OpenStudy (anonymous):

do i need to find the constant of first inegral before integrating it again?

OpenStudy (dumbcow):

yes, thats where the s'(0)=200 comes in plug in 0 for t and set equal to 200 to solve for contant

OpenStudy (anonymous):

okay, thankyou so much

OpenStudy (dumbcow):

your welcome i get s(t) = cos(4t+pi) +200t +1

OpenStudy (anonymous):

i get s(t)=16cos(4t+pi)+200t+16...:/

OpenStudy (dumbcow):

when integrating you should be dividing coefficient by 4 --> u = 4t+pi du = 4 dt dt = du/4

OpenStudy (anonymous):

oh oh oh.. i totally forgot that, i wasnt even substituting. thankyou soo much.. ur a lifesaver

OpenStudy (dumbcow):

:)

OpenStudy (anonymous):

:D